OpenStudy (anonymous):

28c^2-37c-21 factor by grouping

OpenStudy (shubhamsrg):

why dont you try the method which i told earlier.? this ques is no different

OpenStudy (anonymous):

because im not completely sure that i know how to work these out sorry

OpenStudy (shubhamsrg):

inspite of trying to make us do your homework,,you should yourself try it..thats how you'll be comfortable with the subject i'll give a hint though -21 * 28 = -7*3 * 7*4 = -49 * 12 also -49 + 12 =-37

OpenStudy (anonymous):

im not trying to make people do my work im trying to get a better understanding of how to work my problems being that i take this class online its kinda of hard to get the full understanding of how to work out a problem... i do thank you for your help though BTW these problems are not the homework problems they are just similar so that i can understand how to do them....

OpenStudy (shubhamsrg):

hmmn..getting furious little lady ? anyways,, can you do the problems now? these are just some pathetic problems and by practicing just 2-3 ques,,you should have command over it..

OpenStudy (anonymous):

thanks for your help smh its sad you feel that way espec since you are trying to help but just because you understand it right away doesnt mean that everyone will... thanks again ill hire a tutor instead...

OpenStudy (shubhamsrg):

that should be better(for me atleast) // god help the tutor!! amen!! ;)

OpenStudy (shubhamsrg):

break b as m+n where mn =ac.. whats the big deal!!

OpenStudy (anonymous):

i can tell your young but again thanks for your help.......

OpenStudy (shubhamsrg):

wow!! you can tell i am young!! amazing!! congrats infact!! hermes would be so proud ;) and as i said before..pleased to help..

OpenStudy (shubhamsrg):

please dont mind my words much,,am just being sarcastic here..

OpenStudy (anonymous):

.........whateva

OpenStudy (callisto):

To my understanding, I don't think that method is ''factor by grouping'' (even it's a better method :| ) For factor by grouping, you need to take out common factor and group the rest of the term. For your question, 28c^2-37c-21 =28c^2 - 49c + 12c -21 = 7c(4c-7) + 3(4c-7) = (7c+3)(4c-7)

OpenStudy (callisto):

*even -> even though

OpenStudy (anonymous):

thanks callisto!!! your great :)

OpenStudy (callisto):

welcome ._.